# Learning Program Behavioral Models from Synthesized Input-Output Pairs

### v1

@@ 4. Evaluation
We evaluated Modelizer by learning the execution behavior of real-world programs. For every PUT, we collected two datasets of input-output pairs by executing the program with synthesized and real-world inputs respectively. For every evaluation subject, Modelizer additionally learned an inverse model  $M^{-1}$ that corresponds to a hypothetical program  $\textnormal{PUT}^{-1}$, which is expected to implement an inverse of the behavior of the original program  PUT.
@@ 4.1. Research Questions
    How accurate are the Modelizer models that mock the behavior of the  PUT? This question is at the heart of Modelizer: How well do the learned models perform in predicting outputs from inputs?
@@
    How accurate are the Modelizer models that are trained on synthesized data in processing real-world data? This question evaluates the quality of the learned models when tested with real-world data.
@@
    How accurate are the inverse models that mock the behavior of  $\textnormal{PUT}^{-1}$? With RQ3, we evaluate an important application of Modelizer, namely its ability to predict inputs from outputs.
@@
    How fast Modelizer can synthesize a training dataset? This question evaluates the efficiency of the input generation pipeline implemented in Modelizer.
@@
    How fast Modelizer can learn program models? This question evaluates the efficiency of Modelizer core functionality, which is model learning.
@@
    How well do the synthesized samples represent the real-world data? With RQ6, we investigate the quality of the synthesized training set.
  7. RQ7.
    How much synthetic data is required for mocking behavior at which accuracy? With RQ7, we investigate the impact of the size of the synthesized training set on the accuracy of learned models.
  8. RQ8.
    How should synthetic data be generated to achieve the highest accuracy of the learned models? With RQ8, we investigate the impact of the synthetic data generation strategies on the accuracy of learned models.
  9. RQ9.
    How does Modelizer learn the hyperparameters for each data pair combination? This question assesses the impact of hyperparameters (Section 2.3) on the performance of Modelizer.
  10. RQ10.
    Should Modelizer re-use learned hyperparameters for inverse modeling? With RQ10, we check the reusability of found hyperparameters (Section 2.3) with forward models for training inverse models.
  11. RQ11.
    What is the effect of placeholders on behavior learning? This question assesses the impact of placeholders (Section 2.1.2) on the performance of Modelizer.
  12. RQ12.
    What is the effect of tokenizers on learning performance? With RQ12, we investigate the impact of tokenizers (Section 2.2) on the performance of Modelizer.
@@ 4.2. Evaluation Subjects
(MacFarlane, 2006) is an open-source library and command-line tool for converting markups from one format to another. While the library natively supports the translation in both directions(Markdown to HTML and HTML to Markdown), we have tested it only with Markdown to HTML translation functionality. We manually extracted a Markdown Grammar from the CommonMark Spec Version 0.29 (MacFarlane, 2019) specification. We used Pandoc’s functionality to convert Markdown to HTML for our case study, first, to validate the feasibility of neural behavior modeling of black-box systems, second, to check the applicability of the same technique for inverse behavior modeling, when a model has to predict the program input given the output. The promising results of our first experiments encouraged us to test the approach with other subjects.
@@
(msticpy) (Hellen, 2019) is a library for InfoSec investigations, data analysis, and hunting in Jupyter Notebooks authoring for Azure Sentinel, which is a cloud-native security information and event management solution offered by Microsoft for intelligent security analytics and threat intelligence. Modelizer was evaluated on the SQL to KQL Conversion feature of MSTIC, which accepts SELECT SQL queries constructed using a subset of ANSI SQL-92 (ISO/IEC 9075:1992, 1992) as input and produces Kusto Query Language (KQL) queries. KQL is a proprietary query language developed by Microsoft, which is used for querying and analyzing data in Microsoft Azure services. The supported subset of SQL-92 specification was encoded as a CFG grammar by us.
@@
(Belotti, 2020) is a Python library for converting mathematical expressions from AsciiMath representation to Mathematical Markup Language (MathML) and representation. The grammar for synthesizing MathML was manually extracted from the MathML 2.0 Specification (Ausbrooks et al., 2003). The training pairs were collected by converting the synthesized MathML formulas first into format, and then the collected formulas were translated once more into AsciiMath format. This way, we managed to collect three different sets of input-output pairs and test the framework’s capabilities in performing the cross-protocol translations in directions that the given PUT does not currently support.
@@ 4.3. Training Setup
For every subject, we generated training sets of different sizes (5k, 10k, 50k, 100k, 250k, 500k, 1M). The input generation starts with the minimal number of non-terminal expansions (MinNonTerminals) set to 10 and maximal (MaxNonTerminals) set to 20. Once configured, the generator has 100,000 attempts to generate unique samples using a fixed configuration. If the generator runs out of attempts the values of MinNonTerminals and MaxNonTerminals parameters are increased by 10 and the generator is restarted. The generation process will be running till the generator produces the required number of unique samples. Also, we configured the data generators to produce a separate set of input samples that had been synthesized with a continuously increasing complexity of the generated samples. It was achieved by implementing the sliding window modification of MinNonTerminals and MaxNonTerminals values. In this case, the generator uses at least four different non-terminal expansion configurations to generate the required number of samples.
Pandoc, msticpy, and py-asciimath subjects were evaluated using real-world data. We evaluated the SQL to KQL translation models on 1,000 samples that were collected from the “sql-create-context” open-access dataset (brianm, 2023). This dataset contains natural language to SQL SELECT query pairs. During testing the msticpy subject, we ensured that it produced a non-empty output for every extracted SQL query extracted from the dataset. For the py-asciimath subject, we also extracted 1000 real-world MathML formulas from the NTCIR-12 dataset (Zanibbi, 2016). For testing the models of the Pandoc program we extracted 12,783 top-level elements (tree-like structures nesting other markdown definitions) from 100 markdown files that were collected from the open-access Markdown dataset (Schmid, 2023) containing readme documentation files of projects hosted at the HuggingFace Hub (Face, 2024). We have not managed to find a dataset containing mathematical expressions written in Python using the math module from the Python Standard Library. Instead, we synthesized 1000 samples using our Python Expression grammar and replaced the placeholders with random values of corresponding data types. Additionally, we evaluated every model on 10,000 synthesized samples that were not included in the training set. The synthesized dataset was generated by fixing the minimal number of non-terminal expansions to 10 and the maximal number of non-terminal expansions to 20.
@@
While the decoder of Modelizer’s translation routines supports a configurable beam size parameter, we collected all of our experiments with a fixed beam size = 1. All models were trained for 10 iterations(epochs) with a context-window size equal to 5000 tokens. In every training configuration, 80% of the synthesized data had been used for model training. After every training iteration, the training quality was validated using the remaining 20% of the training set. Additionally, the learning quality across models was evaluated with 10,000 previously unseen data pairs which were collected using synthesized inputs, that were generated by setting the maximal number of non-terminal expansion to 20. We trained all models using a single NVIDIA Geforce RTX 4090 GPU with 24 GB of Memory. During the model learning phase Modelizer was allocating from 650 MB to 22 GB of memory depending on the configuration of hyperparameters and the sizes of collected vocabularies. Once trained, models can be deployed on any modern computer architecture that is equipped with at least 8 GB of system memory and can execute a Python interpreter version 3.10 or newer. In contrast to real-world programs the learned models do not require any external dependencies or permissions for system-wide installation. The required Python interpreter and the set of Python modules can be pre-configured and shipped with the model instance. A single environment can be used to execute multiple instances of the model in parallel. Because the size of trained models is relatively small, we monitored a low memory footprint when they were running in the inference mode.
@@ 4.4. Evaluation Metrics
(bilingual evaluation understudy) (Papineni et al., 2002) is a commonly used metric for the evaluation of machine translation systems, that calculates the geometric average precision of single, double, triple, and quadruple token overlap cases between reference and hypothesis token sequences multiplied by brevity penalty. The score output is always a number between 0 and 1, with 1 indicating the perfect translation case. The BLEU score was computed on the whole corpus of test cases using the implementation provided by the Python NLTK package (Team, 2007).
@@
stands for Standard Error computed using BLEU score.
@@
(WER) (Morris et al., 2004) is a metric that calculates the accuracy of a model prediction by computing a ratio between the sum of token insertions, deletions, and substitutions required to transform the predicted token sequence into the reference token sequence and the total number of tokens in the reference token sequence. The lower the value the better the model performs with a score of 0 indicating that the model has not made any mistakes during the translation process.
World Information Lost:
(WIL) (Morris et al., 2004) is a metric that computes the percentage of total tokens that were incorrectly predicted by the model between sets of reference token sequences and sets of hypothesis token sequences. The lower the value the better the model performs with a score of 0 indicating that the model has not lost any information during the translation process.
@@
is a metric that reflects the percentage of evaluation samples that were translated correctly by the model. The higher the value, the better the model performs with a score of 100 indicating that the model has not made any mistakes during the translation process. This metric relies on the Levenshtein distance to compare the similarity between the predicted and reference sequences.
@@ 4.5. Results
Pandoc HTML Markdown Inverse 1M 0.7036 0.1554 7.5202 23.74 34.48 93.27 93.58
@@
latexify-py PyExpres. Inverse 100k 0.8747 0.1240 11.0931 10.20 12.21 44.60 56.00
@@
py-asciimath MathML AsciiMath Hypothet. Inverse 5k 0.2963 0.5779 3.6842 54.06 67.69 3.40 8.80
Yellow cells indicate an evaluation run with synthetic test data, blue cells indicate evaluation runs with the real-world data, and purple cells indicate the evaluation runs on the real-world data performed after fine-tuning the baseline model on a few real-world data pairs.
RQ1: Mocking Accuracy
We start with addressing RQ1: How well do Modelizer models predict outputs from inputs? To answer this question, we evaluated our models on sets of 10,000 synthesized test cases per subject that have no intersection with any of the training sets. The results of the evaluation are summarized in Table 3. The evaluation results show that the models trained on the synthetic data can achieve a high prediction accuracy for most subjects from the evaluation set. For most models and subjects, the BLEU score is above 0.98, the Word Information Lost score is below 0.02 and the Exact Match is above 93%.
@@
Models were trained on the various sizes of datasets containing synthesized Markdown to HTML structures. Blue cells indicate models trained on synthetic data generated with a fixed number of non-terminal expansions, and purple cells indicate application of the synthetic data with an increasing number of non-terminal expansions.
@@
Models were trained on the various sizes of datasets containing synthesized Markdown to HTML structures. Blue cells indicate models trained on synthetic data generated with a fixed number of non-terminal expansions, and purple cells indicate application of the synthetic data with an increasing number of non-terminal expansions.
According to the prediction accuracy scores depicted in Table 3 and representative results from both Table 4 and Table 5, the medium size of training sets can already achieve a high prediction accuracy for most subjects from the evaluation set. Using the given grammar definitions, the data generator can produce the vast majority of data feature combinations within the 100,000–250,000 samples range. By further increasing the number of generated samples, while eliminating duplicates, the generator starts to produce a more unnatural combination of features, which tends to negatively affect the model accuracy, makes the learning more costly, and hits the prediction performance.
@@
Modelizer models trained on synthetic data have a high accuracy in predicting outputs from inputs.
RQ2: Mocking Performance with Real Data
@@
To answer this question, we evaluated our models on sets of 1,000 real-world test cases per subject, except Markdown and HTML, where, as already mentioned, we took 100 real-world readme documents and extracted 12,783 top-level elements from them. These test sets have no intersection with any of the train sets. While the models achieve great accuracy when tested with synthetic data, it is noticeably lower when tested with real-world data.
For the input generation, we did not learn the feature distribution from the particular set of test cases to avoid overfitting to the test data and check Modelizer’s capabilities in learning the program behavior while having no additional knowledge of the test subject and the data it processes. Consecutively, the feature distribution between the real-world and synthesized samples did not align well resulting in a significant performance hit. The availability of additional data pairs for the few-shot fine-tuning of the learned models can significantly boost the prediction accuracy and partially compensate for the feature distribution inconsistency at a small cost. After the fine-tuning, we recorded the edit distance value between expected and predicted output to be greater than 1 mostly for a small fraction of test samples. The inspection of these test cases has shown that they contain features that were not part of the grammar that we used to synthesize input/output pairs, thus samples with similar sets of features were not present in the training data. The only way to fix this issue is to populate the grammar definition with new expansion rules. In contrast to the majority of test data, a few test cases contained an unusual number of placeholder tokens of a particular type. More on this will be discussed in the next research question.
Modelizer requires that the training data (and hence, the synthesizing grammar) covers the features found in the testing data.
In this evaluation, all models that had been trained on MathML Formulas used as input or output (MathML to , to MathML, MathML to AsciiMath, AsciiMath to MathML) have shown the lowest prediction accuracy. We treated them as outliers and investigated the cause of failure. The reason for this is the complexity of the MathML language and the lack of the corresponding test-case features in the grammar we used for input generation. The MathML language supports several operators and operands, including the different encodings for the semantically same element. The grammar we used for the input generation covers only a fraction of these features. Due to this fact, the custom tokenizer also fails to parse these special cases. The better mapping of operators and operands during the tokenization will help to resolve this issue. While the MathML grammar was used for synthetic input generation and the py-asciimath library was utilized for sequential conversion of first MathML to and then to AsciiMath during the training dataset formation, the accuracy of AsciiMath to and to AsciiMath models were less impacted by the inconsistency of the vocabularies. The accuracy drop was only caused by the lack of the specific unseen features, while the encoding variations of the same operators got mapped to the same command by the py-asciimath library in the and AsciiMath representations, thus the model’s vocabulary had already contained them. However, as the Table 3 also depicts, the MathML models manage to learn the conversion behavior for the synthesized test cases. We believe that alterations to the grammar and the tokenization process will help to improve the accuracy of the learned models for MathML Formulas.
In the performance evaluation, MathML is an outlier, due to the restricted grammar used for input generation.
RQ3: Inverse Accuracy
We turn to RQ3 and evaluate the accuracy of Modelizer when predicting inputs from outputs. We again refer to the prediction accuracy scores depicted in Table 3 and representative results from both Table 4 and Table 5. In particular, we are interested in HTML to Markdown, KQL to SQL, to PyExpression, and AsciiMath to models. The numbers show that the inverse model achieves a similar Exact Match accuracy, as their counterpart forward conversion models. The BLEU Score values that we recorded for both forward and inverse models are high and have insignificant margins between them. While the models of the Pandoc and msticpy library achieve more than 90% accuracy in both directions, the latexify-py and py-asciimath can produce fully correct results in half of the cases. Between 20% to 35% of predictions require up to 4 edit operations, including token insertions and deletions, to correct the prediction. Currently, Modelizer can automatically detect these failures using the prediction validation mechanism discussed in Section 2.4. We plan to integrate the automatic prediction repair mechanism in our future work.
Learning inverse models is as performant as learning mock models.
RQ4: Dataset Generation Time
Figure 5. Speed of Modelizer training data generation. This graph represents the data generation time distribution for different subjects and training set sizes. The x-axis represents the input types, and the y-axis represents the time in minutes. Every training set size is represented by a different color. For every input type, the generator was producing samples with fixed and increasing complexity (partitioned).

### v2

@@ 4. Evaluation
We evaluated Modelizer by learning the execution behavior of real-world programs. We collected two datasets of input-output pairs for every PUT by executing each program with synthesized and real-world inputs, respectively. For every evaluation subject, Modelizer additionally learned an inverse model  $M^{-1}$ that corresponds to a hypothetical program  $\textnormal{PUT}^{-1}$, which is expected to implement an inverse of the behavior of the original program  PUT. We also equipped our framework with the ability to fine-tune LLMs to capture the behavior of  PUT and its inverse  $\textnormal{PUT}^{-1}$and compared this approach against custom-trained models.
@@ 4.1. Research Questions
    How accurate are the Modelizer models in mocking the behavior of the  PUT and its inverse  $\textnormal{PUT}^{-1}$? This question is at the heart of Modelizer: How well do the learned models perform in predicting outputs from inputs and vice-versa?
@@
    How well do the synthesized samples represent the real-world data and how much synthetic data is required for mocking program behavior? With RQ2, we investigate the quality of the synthesized training set, its correlation to real-world data, and investigate the impact of the size of the synthesized training set on the accuracy of the learned model.
@@
    How efficiently can Modelizer learn program models? This question evaluates Modelizer’s input generation and model learning efficiency.
@@
    How does Modelizer learn the hyperparameters for each data pair combination? This question assesses the impact of hyperparameters (Section 2.3) on the performance of Modelizer. We also check the reusability of found hyperparameters with forward models for training inverse models.
@@
    What is the effect of tokenizer selection on learning performance? With RQ5, we investigate the impact of tokenizers (Section 2.2) on the performance of Modelizer.
@@
    Could LLMs pre-trained on code mock programs? With RQ6, we investigate the ability of LLMs pre-trained on code replicating program behavior using zero-shot evaluation (without task-specific model adaptation) or after additional fine-tuning and compare their performance with custom-trained models.
@@ 4.2. Evaluation Subjects
(MacFarlane, 2006) is an open-source library and command-line tool for converting markups from one format to another. While the library natively supports the translation in both directions(Markdown to HTML and HTML to Markdown), we have tested it only with Markdown to HTML translation functionality. We manually extracted a Markdown Grammar from the CommonMark Spec Version 0.29 (MacFarlane, 2014) specification. We used Pandoc to convert Markdown to HTML in our case study to validate neural behavior modeling of black-box systems and its applicability for inverse behavior modeling, where the model predicts input from output. The promising results of our first experiments encouraged us to test the approach with other subjects.
@@
(msticpy(Hellen, 2019) is a library for InfoSec investigations, data analysis, and hunting in Jupyter Notebooks for Azure Sentinel, which is a cloud-native security information and event management solution offered by Microsoft for intelligent security analytics and threat intelligence. Modelizer was evaluated on MSTIC’s SQL to KQL Conversion feature, which converts SELECT SQL queries, using a subset of ANSI SQL-92 (ISO/IEC 9075:1992, 1992), into Kusto Query Language (KQL) - a proprietary query language developed by Microsoft for analyzing data in Azure Cloud. We encoded the supported SQL-92 subset as a CFG grammar.
@@
(Belotti, 2020) is a Python library for converting AsciiMath expressions to MathML and formats. The MathML grammar was manually extracted from MathML 2.0 Specification (Ausbrooks et al., 2003). Training pairs were gathered by converting synthesized MathML formulas to and then to AsciiMath. This process yielded three sets of input-output pairs, allowing us to test cross-protocol translations in directions unsupported by the original program.
@@ 4.3. Training Setup
For every subject, we generated training sets of different sizes (5k, 10k, 50k, 100k, 250k, 500k, 1M). The input generation starts with the minimal number of non-terminal expansions (MinNonTerminals) set to 10 and maximal (MaxNonTerminals) set to 20. Once configured, the generator has 100,000 attempts to generate unique samples using a fixed configuration. If the generator runs out of attempts, the values of MinNonTerminals and MaxNonTerminals parameters are increased by 10, and the generator is restarted. The generation process runs till the generator produces the required number of unique samples. Also, we configured the generators to produce a separate set of input samples that had been synthesized with an increasing complexity of the generated samples. It was achieved by implementing the sliding window modification of MinNonTerminals and MaxNonTerminals values. In this case, the generator uses at least four different non-terminal expansion configurations to generate the required number of samples. Inputs were generated on an Ubuntu Linux 22.04 workstation equipped with a 24-core AMD Ryzen Threadripper 3960X CPU and 128 GB of memory.
Each model was evaluated on 10,000 unseen synthesized samples generated with a fixed number of non-terminal expansions between 10 and 20. Additionally, we evaluated models using the following datasets with real-world inputs:
Pandoc.:
We evaluated both Pandoc models with real-world data. We extracted 12,783 top-level elements (tree-like structures nesting other Markdown definitions) from 100 Markdown files that were collected from the open-access Markdown dataset (Schmid, 2023) containing README files of projects hosted at the HuggingFace Hub (Face, 2024).
msticpy.:
We tested msticpy models on 1,000 samples from the “sql-create-context” dataset (b mc2, 2023), featuring natural language and SQL SELECT query pairs. We ensured that  PUTproduced non-empty outputs for every selected SQL query.
latexify-py.:
The evaluation of latexify-py was challenging due to the lack of datasets with mathematical expressions written in Python using the math module from the Python Standard Library. Thus, we synthesized 1,000 additional samples using our Python Expression grammar and replaced placeholders with random values of the appropriate type.
py-asciimath.:
The evaluation of py-asciimath involved 1,000 real-world MathML formulas from the NTCIR-12 dataset (Zanibbi, 2016).
@@
While Modelizer’s translation routines support a configurable decoder beam size, we fixed the beam size for our experiments to 1. All models were trained for 10 epochs with a context window of 5000 tokens. In each configuration, 80% of synthesized data was used for training, and the training quality was validated with the remaining 20%. All models were trained on a single NVIDIA Geforce RTX 4090 GPU with 24 GB memory. During learning, Modelizer allocated 650 MB to 22 GB of memory based on hyperparameters and vocabulary size values. All models trained from scratch can be deployed on modern computers with at least 8 GB memory that can execute a Python interpreter version 3.10 or newer. Unlike real-world programs, these models require no external dependencies or system-wide installation permissions. The necessary Python interpreter and modules can be pre-configured and shipped with the model. A single environment can run multiple model instances in parallel due to a lower memory footprint at the inference time.
Additionally, we experimented with querying and fine-tuning two commonly used LLMs pre-trained on code: CodeLlama (Rozière et al., 2024) and CodeGemma (Team et al., 2024). The selected models were instruction tuned by their vendors, meaning that vendors have fine-tuned them on natural language instruction followed by input and expected output pairs. This process improves the model’s general utility and performance in practical applications. Both models have 7 billion trainable parameters, and their weights were additionally quantized to 4bit to reduce memory requirements. Experiments involving the usage of LLMs were performed on the data-center grade NVIDIA A100 Tensor Core GPU with 40GB memory. At first, models were queried using the same real-world test samples from previous experiments. After that, we performed fine-tuning on synthesized data and repeated the evaluation with test samples. We fine-tuned an individual model for 10 epochs for each input-output pair to exclude potential interference between subjects and simplify comparison with custom-trained models. During fine-tuning, we reused the synthesized training sets with 10,000 samples generated using the MinNonTerminals) set to 10 and MaxNonTerminals set to 20.
@@ 4.4. Evaluation Metrics
(bilingual evaluation understudy) (Papineni et al., 2002) is a commonly used metric for the evaluation of machine translation systems that calculates the geometric average precision of single, double, triple, and quadruple token overlap cases between reference and hypothesis token sequences multiplied by brevity penalty. The score output is always a number between 0 and 1, with 1 indicating the perfect translation case. The BLEU score was computed on the whole corpus of test cases using the implementation provided by the Python NLTK package (Team, 2007).
@@
stands for the Standard Error computed using the BLEU score.
@@
(WER) (Morris et al., 2004) measures prediction accuracy as a ratio of token insertions, deletions, and substitutions needed to convert the predicted token sequence into the reference sequence and the total reference token count. A lower WER indicates better performance, with a score of 0 showing no translation mistakes.
Word Information Lost:
(WIL) (Morris et al., 2004) measures the percentage of total tokens incorrectly predicted between reference and hypothesis token sequences. The lower the value, the better the model performs, with a score of 0 indicating that the model has not lost any information during the translation process.
@@
measures the percentage of evaluation samples correctly translated by the model. A higher value indicates better performance, with a score of 100 showing no translation errors. This metric uses Levenshtein distance to assess the similarity between predicted and reference sequences.
@@ 4.5. Results
Pandoc HTML Markdown Backward 1M 0.7036 0.1554 7.5202 23.74 34.48 93.27 93.58
@@
latexify-py PyExpres. Backward 100k 0.8747 0.1240 11.0931 10.20 12.21 44.60 56.00
@@
py-asciimath MathML AsciiMath Hypothet. Backward 5k 0.2963 0.5779 3.6842 54.06 67.69 3.40 8.80
Yellow cells = evaluation with synthetic data; Blue cells = evaluation with real-world data;
purple cells = evaluation with real-world data after few-shot fine-tuning.
RQ1: How accurate are the Modelizer models in mocking the behavior of the  PUT and its inverse  $\textnormal{PUT}^{-1}$?
We start with addressing RQ1: How well do Modelizer models predict outputs from inputs? To answer this question, we evaluated our models on sets of 10,000 synthesized test cases per subject that have no intersection with any of the training sets. The results of the evaluation are summarized in Table 3. They show that the models trained on the synthetic data can achieve a high prediction accuracy for most subjects from the evaluation set. For most models and subjects, the BLEU score is above 0.98, the Word Information Lost score is below 0.02 and the Exact Match is above 93%.
Modelizer models mock the program behavior without making errors in more than 93% of cases.
Next, we evaluated the accuracy of Modelizer when predicting inputs from outputs. The prediction accuracy scores of HTML to Markdown, KQL to SQL, to PyExpression, and AsciiMath to models depicted in Table 3 show that the inverse model achieves a similar Exact Match accuracy as their counterpart forward conversion models. The recorded BLEU Score values for both forward and inverse models are high and have insignificant margins between them. While the models of the Pandoc and msticpy library achieve more than 90% accuracy in both directions, the latexify-py and py-asciimath can produce fully correct results in half of the cases. Between 20% to 35% of predictions require up to 4 edit operations, including token insertions and deletions, to correct the prediction. Currently, Modelizer can automatically detect these failures using the prediction validation mechanism discussed in Section 2.4. We plan to integrate the automatic prediction repair mechanism to tackle inaccurate prediction cases in our future work.
@@
Blue cells = fixed number of expansions; purple cells = increasing number of expansions.
@@
Blue cells = fixed number of expansions; purple cells = increasing number of expansions.
@@
Modelizer models trained on synthetic data are also accurate in predicting inputs from outputs.
According to the representative results for Pandoc models depicted in Table 4 and Table 5, models trained on the medium training set size can already achieve a high prediction accuracy. Using the input generation configuration specified in Section 4.3, the data generator can produce the vast majority of input feature combinations within the 100,000–250,000 samples range. By further increasing the number of generated samples while eliminating duplicates, the generator starts to produce more unnatural input feature combinations, which tends to negatively affect the model accuracy by making the learning more costly and hitting the prediction performance.
Medium datasets with 100,000 to 250,000 synthesized samples suffice to obtain good Modelizer models.
RQ2: How well do the synthesized samples represent the real-world data and how much synthetic data is required for mocking program behavior?
@@
To answer this question, we evaluated our models on sets of 1,000 real-world test cases per subject, except Markdown and HTML, where, as already mentioned, we took 100 real-world readme documents and extracted 12,783 top-level elements (basic blocks) from them. These test sets do not intersect with any training sets. According to Table 3, while all models achieved high accuracy with synthetic test data, they performed noticeably worse when tested with real data.
A probable cause for this effect is that synthesized and real-world test sets have different distributions of input-output features. In input generation, we intentionally avoided learning feature distribution from specific test cases to prevent overfitting and to assess Modelizer’s ability to learn program behavior without extra knowledge of the test subject or processed data. As a result, the feature distribution between real-world and synthesized samples did not align, causing a significant performance drop. This negative effect can be fixed with few-shot fine-tuning (Brown et al., 2020; Schick and Schütze, 2021), a widely used technique for the adaptation of pre-trained models to domain-specific tasks at a small cost. We applied it to models trained on synthetic data, which helped to mitigate feature distribution issues and enhanced prediction accuracy. Performance improved significantly after fine-tuning each model with 10 new randomly selected input-output pairs from source datasets. Purple cells in Table 3 reflect the performance numbers after few-shot fine-tuning.
Figure 5. Markdown Sequence Length Frequency Distribution. The x-axis represents the length of different token sequences, and the y-axis denotes the frequency count of each token length. Every data split is represented by a different color.